OpenStudy (anonymous):

Explain why it is necessary to write "+C" when finding an antiderivative i know the stands for (C for the constant of integration) but i don't know why its nesessary

OpenStudy (abb0t):

the "C" is to cover all possible solutions.

OpenStudy (dumbcow):

its because the derivative of a constant is zero. so many similar functions could all have same derivative f(x) = x+2 , g(x) = x +3 f'(x) = g'(x) =1 thus \[\int\limits_{}^{} 1 dx = x + C\]

OpenStudy (anonymous):

thank you so much (:

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):